A rectangle is 5 times as lon as it is wide. If the area is 125 square feet, find the perimeter.
Answer by checkley77(12844)

L=5W
LW=125
5W*W-125
5W^2=125
W^2=125/5
W^2=25
W=SQRT25
W=5 ANS. FOR THE WIDTH.
L=5*5=25 ANS FOR THE LENGTH.
PROOF:
5*25=125
125=125